# NVM-in-Cache: Repurposing Commodity 6T SRAM Cache into NVM Analog Processing-in-Memory Engine using a Novel Compute-on-Powerline Scheme

### v1

@@ V-E Accuracy Analysis
To emulate hardware-induced variations, we injected Gaussian noise with variable standard deviations estimated from Monte Carlo simulations into the ADC output. This perturbation was applied after the inverse mapping, with its amplitude scaled to the dynamic range of the activation outputs, ensuring that the simulated quantization faithfully captured target hardware conditions. The experimental outcomes are summarized in Table II. The baseline ResNet-18 model achieved an accuracy of 91.84%. After fine-tuning for 200 epochs with an SGD optimizer (learning rate 0.001, cosine annealing schedule), enabling non-linear ADC quantization without noise reduced accuracy to 91.55%, and the inclusion of scaled Gaussian noise caused a further drop to 91.27%. In contrast, without fine-tuning, the model achieved only $\sim$}77% accuracy on CIFAR-10, underscoring the necessity of task-aware adaptation when deploying under realistic ADC non-idealities.
@@ V-F Comparison with previous works
Table I provides a detailed comparison of the proposed 6T-2R NV-SRAM based PIM architecture with state-of-the-art PIM designs. Prior works have primarily relied on time/charge-domain SRAM [35], [38], [39] or 1T1R RRAM/MRAM arrays [37, 40] to demonstrate in-memory computing. In contrast, our work integrates RRAM devices within a conventional SRAM framework, thereby preserving the native bit-cell area and compatibility with standard memory arrays. This makes the proposed bitcell a drop-in replacement for existing SRAM-based cache. In terms of throughput and energy efficiency, the proposed design achieves $0.4$ TOPS and $491.78$ TOPS/W when adjusted to 1-bit precision, which is comparable to prior implementations of PIM based on CMOS and ReRAM. The normalization to 1-bit precision follows standard convention in CIM/PIM benchmarking, ensuring fair comparison across different input/output precisions. It should be noted that the latency bottleneck in our proposed scheme originates primarily from the 6-bit SAR ADC, which operates at $50~\text{MHz}$. While sufficient for functional demonstration, higher-speed ADCs or parallel conversion can significantly improve latency, which would directly increase the achievable throughput (TOPS). Moreover, the ADC also contributes a substantial portion of the reported area. Future solutions, such as 3D integration or ADC sharing across multiple columns or cache banks, can greatly reduce this overhead. Since the proposed 6T-2R bit-cell occupies the same area as a conventional SRAM cell, the overall bit-cell area density remains unaffected, and improvements in ADC design will directly translate into both higher compute density (TOPS/mm2) and higher throughput efficiency (TOPS/W). Taken together, these results highlight that while earlier works have shown CIM functionality using charge-domain SRAM or RRAM/MRAM crossbars, our approach uniquely demonstrates SRAM-compatible PIM with embedded RRAM, combining the benefits of area neutrality, non-volatility, and competitive throughput/energy efficiency. This makes the proposed design well-suited as a scalable drop-in replacement for conventional SRAM macros in future cache and memory hierarchies.
@@ VI Conclusion
This paper presented an NVM-in-Cache architecture based on a 6T-2R hybrid bit-cell that integrates resistive RAM (RRAM) devices within a conventional SRAM framework. The proposed design doubles the overall memory capacity and enables NVM in-memory computation without altering the native cell footprint, making it a drop-in replacement for standard SRAM cache. By retaining cache data during computation, the architecture eliminates costly cache flush and reload operations while simultaneously enhancing storage density. Circuit-level simulations in 22 nm FDSOI confirm robust resistive switching, with fast 4 ns programming and a high HRS/LRS ratio that ensures reliable binary storage. Furthermore, integration with a calibrated 6-bit SAR ADC achieves full code utilization, maintaining monotonicity and efficient digital mapping of analog weights. System-level evaluations show a throughput of $0.4$ TOPS and an energy efficiency of $491.78$ TOPS/W, demonstrating performance. For 128 row-parallel operations, the CIFAR-10 classification achieves an accuracy of 91.27%. While the ADC currently dominates both latency and area, future improvements such as 3D stacking and ADC sharing across columns can further boost throughput and compute density. This work does not preclude implementing purely SRAM-based PIM schemes (without NVM); rather, it can be used in conjunction with such prior approaches, effectively doubling the memory capacity and expanding the opportunities to realize IMC within conventional 6T cells. Overall, the proposed 6T-2R NV-SRAM provides a scalable and energy-efficient solution for next-generation cache hierarchies and AI accelerators.

### v2

@@ V-E Accuracy Analysis
To emulate hardware-induced variations, we injected Gaussian noise with variable standard deviations estimated from Monte Carlo simulations into the ADC output. This perturbation was applied after the inverse mapping, with its amplitude scaled to the dynamic range of the activation outputs, ensuring that the simulated quantization faithfully captured target hardware conditions. The experimental outcomes are summarized in Table II. The baseline ResNet-18 model achieved an accuracy of 92.88%. After fine-tuning for 200 epochs with an SGD optimizer (learning rate 0.001, cosine annealing schedule), enabling non-linear ADC quantization without noise reduced accuracy to 92.00%, and the inclusion of scaled Gaussian noise caused a further drop to 91.76%. In contrast, without fine-tuning, the model achieved only $\sim$}77% accuracy on CIFAR-10, underscoring the necessity of task-aware adaptation when deploying under realistic ADC non-idealities.
@@ V-F Comparison with previous works
Table I provides a detailed comparison of the proposed 6T-2R NV-SRAM based PIM architecture with state-of-the-art PIM designs. Prior works have primarily relied on time/charge-domain SRAM [TCAS2], [chen2023samba], [song202328] or 1T1R RRAM/MRAM arrays [isscc22, nverma] to demonstrate in-memory computing. In contrast, our work integrates RRAM devices within a conventional SRAM framework, thereby preserving the native bit-cell area and compatibility with standard memory arrays. This makes the proposed bitcell a drop-in replacement for existing SRAM-based cache. In terms of throughput and energy efficiency, the proposed design achieves $0.4$ TOPS and $452.34$ TOPS/W when adjusted to 1-bit precision, which is comparable to prior implementations of PIM based on CMOS and ReRAM. The normalization to 1-bit precision follows standard convention in CIM/PIM benchmarking, ensuring fair comparison across different input/output precisions. It should be noted that the latency bottleneck in our proposed scheme originates primarily from the 6-bit SAR ADC, which operates at $50~\text{MHz}$. While sufficient for functional demonstration, higher-speed ADCs or parallel conversion can significantly improve latency, which would directly increase the achievable throughput (TOPS). Moreover, the ADC also contributes a substantial portion of the reported area. Future solutions, such as 3D integration or ADC sharing across multiple columns or cache banks, can greatly reduce this overhead. Since the proposed 6T-2R bit-cell occupies the same area as a conventional SRAM cell, the overall bit-cell area density remains unaffected, and improvements in ADC design will directly translate into both higher compute density (TOPS/mm2) and higher throughput efficiency (TOPS/W). Taken together, these results highlight that while earlier works have shown CIM functionality using charge-domain SRAM or RRAM/MRAM crossbars, our approach uniquely demonstrates SRAM-compatible PIM with embedded RRAM, combining the benefits of area neutrality, non-volatility, and competitive throughput/energy efficiency. This makes the proposed design well-suited as a scalable drop-in replacement for conventional SRAM macros in future cache and memory hierarchies.
@@ VI Conclusion
This paper presented an NVM-in-Cache architecture based on a 6T-2R hybrid bit-cell that integrates resistive RAM (RRAM) devices within a conventional SRAM framework. The proposed design doubles the overall memory capacity and enables NVM in-memory computation without altering the native cell footprint, making it a drop-in replacement for standard SRAM cache. By retaining cache data during computation, the architecture eliminates costly cache flush and reload operations while simultaneously enhancing storage density. Circuit-level simulations in 22 nm FDSOI confirm robust resistive switching, with fast 4 ns programming and a high HRS/LRS ratio that ensures reliable binary storage. Furthermore, integration with a calibrated 6-bit SAR ADC achieves full code utilization, maintaining monotonicity and efficient digital mapping of analog weights. System-level evaluations show a throughput of $0.4$ TOPS and an energy efficiency of $452.34$ TOPS/W, demonstrating performance. For 128 row-parallel operations, the CIFAR-10 classification achieves an accuracy of 91.27%. While the ADC currently dominates both latency and area, future improvements such as 3D stacking and ADC sharing across columns can further boost throughput and compute density. This work does not preclude implementing purely SRAM-based PIM schemes (without NVM); rather, it can be used in conjunction with such prior approaches, effectively doubling the memory capacity and expanding the opportunities to realize IMC within conventional 6T cells. Overall, the proposed 6T-2R NV-SRAM provides a scalable and energy-efficient solution for next-generation cache hierarchies and AI accelerators.